### v2

# GEIC: Universal and Multilingual Named Entity Recognition with Large Language Models

Abstract
Large Language Models (LLMs) have supplanted traditional methods in numerous natural language processing tasks. Nonetheless, in Named Entity Recognition (NER), existing LLM-based methods underperform compared to baselines and require significantly more computational resources, limiting their application. In this paper, we introduce the task of generation-based extraction and in-context classification (GEIC), designed to leverage LLMs’ prior knowledge and self-attention mechanisms for NER tasks. We then propose CascadeNER, a universal and multilingual GEIC framework for few-shot and zero-shot NER. CascadeNER employs model cascading to utilize two small-parameter LLMs to extract and classify independently, reducing resource consumption while enhancing accuracy. We also introduce AnythingNER, the first NER dataset specifically designed for LLMs, including 8 languages, 155 entity types and a novel dynamic categorization system. Experiments show that CascadeNER achieves state-of-the-art performance on low-resource and fine-grained scenarios, including CrossNER and FewNERD. Our work is openly accessible.
GEIC: Universal and Multilingual Named Entity Recognition
with Large Language Models
Hanjun Luothanks: Zhejiang University, hanjun.21@intl.zju.edu.cn 11footnotemark: 1, Yingbin Jinthanks: Zhejiang University, 3210104412@zju.edu.cn 11footnotemark: 1, Xuecheng Liuthanks: Zhejiang University, xuecheng.21@intl.zju.edu.cn 11footnotemark: 1, Tong Shangthanks: Southeast University, 213211756@seu.edu.cn 22footnotemark: 2, Ruizhe Chenthanks: Zhejiang University, ruizhec.21@intl.zju.edu.cn 11footnotemark: 1, Zuozhu Liuthanks: Zhejiang University, zuozhuliu@intl.zju.edu.cn, Corresponding author 11footnotemark: 1
1 Introduction
The rapid advancement of Large Language Models (LLMs) has revolutionized natural language processing (NLP) (Zhao et al., 2023; Naveed et al., 2023). Recent LLMs, such as GPT-4 (Achiam et al., 2023) and Claude-3 Anthropic (2023), with their powerful generalization and attention mechanisms Vaswani et al. (2017), excel not only in text generation tasks like Text Summarization Wilson et al. (2024); Jin et al. (2024), Question Answering Tan et al. (2023), and Machine Translation Vilar et al. (2022); Moslem et al. (2023), but also in discriminative tasks, including Named Entity Extraction Sancheti et al. (2024) and Text Classification Gasparetto et al. (2022).
Nevertheless, unlike other discriminative tasks, existing LLM-based approaches Shao et al. (2023); Keloth et al. (2024); Li and Zhang (2023) struggle to replace supervised methods in Named Entity Recognition (NER). NER is typically framed as a sequence labeling task, where the model assigns entity labels to each token in a sentence. This structure emphasizes sequential dependencies, which contrasts with LLMs’ focus on global attention mechanisms. Although existing methods leverage generalization and prior knowledge of LLMs to outperform supervised baselines in low-resource scenarios, they often fall short significantly in standard scenarios or complex entity classifications. Moreover, their reliance on commercial models, typically GPT, introduces significant challenges related to cost, data security, and transparency, limiting their broader applicability.
To address these issues, we introduce Generation-based Extraction and In-context Classification (GEIC), a novel task designed to harness LLMs’ attention mechanisms and prior knowledge for NER. As illustrated in Figure 1, unlike traditional sequence labeling methods, GEIC redefines NER as two sequential in-context text generation sub-tasks, extraction and classification, which decreases the task complexity and enhances the capability to capture in-context dependencies. To achieve GEIC, we propose CascadeNER, a universal and multilingual GEIC framework. CascadeNER employs model cascading Varshney and Baral (2022), assigning the each sub-tasks to corresponding fine-tuned small-parameter LLMs, referred to as Small Language Models (SLMs). This method reduces computational resource consumption while maintaining high performance, enabling efficient NER with SLMs. Leveraging the prior knowledge, CascadeNER can complete NER tasks in multiple languages in both few-shot and zero-shot scenarios. Extensive experiments demonstrate that CascadeNER achieves state-of-the-art (SOTA) performance on low-resource and fine-grained datasets. Considering the lack of datasets designed for LLM-based NER, we collect AnythingNER for CascadeNER, a multilingual and fine-grained NER dataset, featuring 8 major languages, 8 coarse-grained types, 31 medium-grained types, 155 fine-grained types, and a novel dynamic categorization system. As the first dataset optimized for LLMs’ generalization and few-shot/zero-shot learning in NER, AnythingNER provides an unprecedented level of detail and coverage. Moreover, this work offers the first comprehensive comparison and analysis of existing LLM-based NER methods, with a emphasis on multilingual and fine-grained scenarios.
Figure 1: Red arrows indicate that the data is used for training. Traditional methods sequentially label each token according to predefined categories from the train set. GEIC does not differentiate between entity types when extracting but introduces a classification sub-task, thereby improving its generalization. The categories can be changed without re-training, enabling zero-shot reasoning.
Our core contributions are summarized as follows:
  • We introduce generation-based extraction and in-context classification (GEIC), a novel task to replace traditional NER tasks in LLMs, emphasizing global attention and generalization
  • We propose CascadeNER, the first universal and multilingual NER framework with SLMs, which supports both few-shot and zero-shot scenarios and achieves SOTA performance on low-resource and fine-grained datasets
  • We build AnythingNER, the first multilingual and fine-grained datasets designed for NER with LLMs, especially GEIC, with a novel dynamic categorization system
  • We conduct comprehensive comparison of existing LLM-based NER methods, providing future directions of LLM-based NER
2 Related Works
2.1 Named Entity Recognition
Named Entity Recognition (NER) is the task of identifying named entities in text and classifying it into predefined categories. Supervised methods, such as BiLSTM Yu et al. (2020) and BERT-MRC Li et al. (2019a), currently dominate this task. They generally rely on large amounts of train data to achieve strong performance, which limits their application in low-resource scenarios. With the advancement of LLMs, leveraging their few-shot and zero-shot learning capabilities to address this. Xie et al. (2023) explore the application of GPT-3 Brown (2020) for NER, but its performance is significantly lower than supervised methods’. GPT-NER Wang et al. (2023) improve GPT-3’s NER capabilities by re-framing the task as single-entity labeling through prompt engineering. While GPT-NER enhanced performance, it still lags behind supervised methods. PromptNER Ashok and Lipton (2023) achieves better results with GPT-4, yet failed to deliver a robust zero-shot solution. Furthermore, several studies investigate LLM-based NER in domain-specific tasks Li and Zhang (2023); Shao et al. (2023); Keloth et al. (2024). These studies, however, all rely on commercial models, increasing costs and limiting flexibility. Existing results highlight both the potential of LLMs in NER and the pressing need for more adaptable and cost-effective solutions.
2.2 Small Language Models
While commercial LLMs undergoes rapid development, small-parameter and open-source LLMs, known as small language models (SLMs), also attracts significant attention Touvron et al. (2023); Hu et al. (2024a); Zhang et al. (2024); GLM et al. (2024); Bai et al. (2023). These models employ techniques such as quantization Han et al. (2015) and distillation Hinton (2015) to reduce their parameters to between 1.5B and 7B,offering advantages like easier fine-tuning and lower computational resource consumption, albeit at the cost of some performance degradation in complex tasks. This trade-off have led to the widely use of SLMs like LLaMA-2-7B in both academia and industry Clemmer et al. (2024); Gade et al. (2023); Luo et al. (2024). However, NER, as a relatively complex task, places high demands on model performance. Existing LLM-based methods fail to effectively decomposed the task, which exacerbates the performance limitations of SLMs in handling complex tasks. Consequently, SLMs have not been effectively applied to NER.
2.3 NER Dataset
A considerable number of NER datasets in various domains have been developed due to the significance of NER Tjong Kim Sang (2002); Kim et al. (2003); Doddington et al. (2004); Walker et al. (2006); Weischedel et al. (2011); Pradhan et al. (2013); Derczynski et al. (2017). However, these existing datasets exhibit several limitations, making them unsuitable for the GEIC task. Most previous multilingual NER datasets adopt coarse-grained classification. For example, PAN-X Pan et al. (2017) includes only three types. Such datasets no longer meet the fine-grained requirements of contemporary flat NER applications. Even existing fine-grained datasets demonstrate clear limitations in category coverage and granularity, falling short of being truly "universal." For instance, FewNERD, despite having 66 entity types, suffers from highly imbalanced data distribution, which affects its reliability for evaluating few-shot learning capabilities. Furthermore, current datasets fail to adequately address the generalization and few-shot/zero-shot learning capabilities of LLMs, hindering the comprehensive training and evaluation of LLM-based NER methods.
3 GEIC
3.1 Definition
Typical NER methods regards NER as a sequence labeling task, where each word is assigned a label, usually BIO representation, and use models to predict entity labels by maximizing the conditional probability of the label sequence:
$P(Y|X)=\prod_{t=1}^{T}P(y_{t}|X,y_{1:t-1})$ (1)
where $X=\{x_{1},\dots,x_{T}\}$ is the input sequence, and $Y=\{y_{1},\dots,y_{T}\}$ is the corresponding label sequence. Such methods focus on local dependencies, making them less effective at capturing sentence-level dependencies. Their reliance on annotated data also limits their adaptability, especially in low-resource scenarios.
Figure 2: Use the sentence "Nothstein beat Buran from Czech Republic" and the multi-granularity categories of AnythingNER as the example. The extractor and classifier are the two different SLMs used in CascadeNER. Azure boxes represent the specific type list for the input of the classifier. Blue boxes represent the sentence input.
Generation-based Extraction and In-context Classification (GEIC), unlike traditional NER tasks, divides NER into two sequential, independently executed, generation-based sub-tasks. In the first sub-task, extraction, the model generate a sentence where all named entities are marked with identifiers and individually re-embeds each entity back into its context, resulting in sentences with identifiers at the number of entities. In the second sub-task, classification, the model receives sentences with identifiers and a list of entity types, and label one entity at a time. When handling multi-granularity data, GEIC allows a hierarchical approach, firstly performing coarse-grained classification and then fine-grained classification based on the identified coarse-grained type and its subcategories.
The unique design of GEIC provides several key advantages. First, GEIC decomposes the NER task into two simpler and independently executed sub-tasks, reducing the complexity of each while enabling the model to better handle diverse and intricate language patterns. This separation allows the sub-tasks to be managed by different models, offering greater flexibility for aggressive optimization strategies that might hinder comprehensive performance. Second, both sub-tasks in GEIC are generation-based and thus well-suited to LLMs. This allows GEIC to leverage the prior knowledge, generalization, and reasoning capabilities of LLMs better, leading to more accurate results, particularly in low-resource scenarios. Third, by performing extraction and classification based on the entire sentence, GEIC aligns with LLMs’ global attention mechanisms, enhancing the relization of contextual relationships. Ultimately, GEIC supports progressive multi-granularity classification, significantly improving performance in handling nuanced and accurate entity classification, making it ideal for the fine-grained requirements of NER applications.
3.2 CascadeNER
To implement GEIC tasks, we propose CascadeNER, a novel framework based on LLMs. A simplified pipeline of CascadeNER with an example is illustrated in Figure 2. To optimize performance while reducing computational resources, CascadeNER employs model cascading, where the extraction and classification sub-tasks are handled separately by two specialized, fine-tuned LLMs, typically SLMs. This structure allows each model to focus on its specific sub-task, maximizing performance on simpler, more specialized tasks and fully aligning with the concept of GEIC.
Upon receiving the input sentence, CascadeNER first processes the sentence by the extractor to mark all named entities with identifiers, and re-embeds each entity back, resulting in sentences with identifiers around the named entities. These sentences are then individually fed into the classifier, which classifies each entity based on the context and the input category list. For multi-granularity categories, CascadeNER employs progressive classification, moving from coarse to fine categories, enhancing its accuracy in handling complex classifications. The detailed steps of extraction and classification are discussed in the following sections.
3.2.1 Extraction
In the extraction sub-task, we utilize a generation-based extraction method, where special tokens "##" are used to surround any entities identified in the sentence, regardless of the number of entities or their types For example:
Query: Kobe was in NBA
Response: ##Kobe## was in ##NBA##
This method, compared to conventional sequential extraction, avoids requiring LLMs to perform text alignment, thus reducing task complexity. In contrast to similar methods Wang et al. (2023); Hu et al. (2024b), CascadeNER’s query contains only the sentence, without any task descriptions, demonstrations, or category information. The response exclusively uses "##" as the identifiers, and all entities are extracted without specifying categories. CascadeNER achieves low-cost NER by using simple prompts and better generalization by treating all entities uniformly. Moreover, extracting in the sentence-level leverages LLMs’ self-attention mechanisms for global understanding, improving accuracy in low-context scenarios. A detailed comparison with existing methods and further advantages of our method are shown in Appendix D.
After conducting extensive experiments, we find that the extractor’s precision consistently exceeds recal, regardless of the model or dataset, indicating that while correct entities are effectively identified, there is a tendency for under-detection. To mitigate this issue, we introduce a union strategy in result fusion Ganaie et al. (2022), allowing multiple extraction for one sentence and taking the union of the results to maximize recall. For cases of entity nesting, where different extraction rounds produce overlapping or nested entities, we apply a length-first strategy, retaining the longer entity, as longer entities generally carry more semantic meaning Nguyen and Cao (2010). For example, "Boston University" is semantically more accurate than "Boston" in the context of "She studies in Boston University". The formula of our strategy is shown below:
$E_{\text{final}}=\bigcup_{i=1}^{n}\left\{\arg\max_{e\in E_{i}}\text{length}(e)\right\}$ (2)
where $E_{i}$ is the set of extracted entities from the $i$-th extraction, $n$ is the number of extraction rounds, $E_{\text{overlap}}$ is the set of overlapping or nested entities, and $\text{length}(e)$ is the length of entity $e$. The effects of the number of extraction repetitions and other details are provided in Appendix C.1.
3.2.2 In-context Classification
In the classification sub-task, we employ a generation-based in-context classification method, where we input the categories and the sentence with one entity surrounded by "##", and require the classifier to generate the label for that entity. This method re-embeds the entity into the sentence for classification, which utilizes the self-attention architecture of LLMs for contextual information and improves accuracy compared to entity-level classification. Figure 3 is an example:
Figure 3: Examples of classification.
For multi-granularity data, we apply hierarchical classification, ensuring the accuracy of classification for complex categories. After obtaining the coarse-grained result, CascadeNER use the result to index the corresponding sub-categories and classify again, continuing this process until no further classification is possible:
$L_{i}^{\text{fine}}=f_{\text{fine-classify}}(L_{i}^{\text{coarse}},\text{subcategories})$ (3)
where $L_{i}^{\text{fine}}$ is the fine-grained label, $L_{i}^{\text{coarse}}$ is the generated coarse-grained label, and subcategories are the subcategories under the coarse-grained.
In zero-shot scenarios, we use a slightly different prompt. Due to significant differences in entity classifications across datasets, some named entities in one dataset may be overlooked in others. We append the query with If none of them applied, return unknown to handle situations where the extracted entity cannot be classified into the provided categories, enhancing CascadeNER’s generalization across datasets.
3.2.3 Few-shot Data Sampling
In existing datasets, only CrossNER Liu et al. (2021), designed for low-resource scenarios, and FewNERD Ding et al. (2021), designed for few-shot scenarios, meet our requirements for evaluating CascadeNER in few-shot scenarios. However, relying solely on them is insufficient for comprehensively evaluating CascadeNER, particularly its multilingual NER performance.
To address this, we develop a sampling algorithm to construct datasets for few-shot evaluation. Considering that basic random sampling cannot ensure a balanced category distribution, we employ a stratified sampling algorithm, which divides the dataset into strata based on the labels. Each stratum corresponds to a distinct entity type, and we ensure an relatively equal number of samples per category by drawing from these strata, thereby maintaining balance across categories in the results. The size for each stratum is calculated with the formula:
$s_{i}=\min\left(\left\lfloor\frac{S}{m}\right\rfloor,n_{i}\right)$ (4)
where $N$ is the total number of labels in the dataset, $S$ is the total sample size, $n_{i}$ is the total number of labels with value $i$, $m$ is the number of categories, and $s_{i}$ is the number of labels from stratum $i$.
4 AnythingNER Dataset
Due to the tremendous parameters and the self-attention mechanism, LLMs demonstrate uniquely high performance in low-resource scenarios for NER tasks. However, there are no existing NER datasets optimized for the characteristics of LLMs, limiting the comprehensive training and evaluation of LLMs in these tasks and making it difficult to effectively evaluate GEIC. Moreover, there is also a lack of multilingual and fine-grained datasets.
Figure 4: The 8 coarse-grained categories, 31 medium-grained categories of AnythingNER. More detailed categories are provided in Appendix G
We thus constructed AnythingNER, the first NER dataset specifically designed for LLM-based NER, particularly GEIC tasks. AnythingNER spans 8 major languages: English, Chinese, Spanish, French, German, Japanese, Korean, and Russian. In terms of classification, it is the first NER dataset to feature three-level granularity categorization, encompassing 8 coarse-grained types, 31 medium-grained types, and 155 fine-grained types, as illustrated in Figure 4.
Compared to existing datasets, AnythingNER offers several key advantages. It adopts a three-level granularity categorization with 155 categories, ensures comprehensive coverage across diverse domains, including rare entities and emerging fields like technology products, which are not adequately considered in earlier datasets. It is specifically formatted to be compatible with GEIC tasks and supports 8 major languages to leverage LLMs’ multilingual capabilities. Additionally, the dataset ensures balanced category distribution for few-shot scenarios and applies a novel dynamic categorization for zero-shot scenarios, making it highly suitable for current research focuses.
Figure 5: Example of GEIC format.
AnythingNER provides data in three formats: the standard BIO format, a novel format for GEIC tasks termed GEIC format, and the dynamic categorized GEIC format. The BIO format follows the same structure as classic datasets like CoNLL2003 Tjong Kim Sang and De Meulder (2003). The structure of the GEIC format is shown in Figure 5. This format is easy to index and includes language information, facilitating data processing, especially in multilingual scenarios. The standard GEIC version is designed for few-shot scenarios, while the dynamic categorized version is for zero-shot scenarios. The core of our dynamic categorization is dynamically adjusting and simplifying the entity type list provided to the model for improving the model’s generalization and reducing the risk of overfitting. Detailed implementation and advantages are provided in Appendix A.
All the corpora in AnythingNER are sourced from Wikipedia contributors (2024) and manually annotated by our team. A statement regarding the annotators and ethical considerations can be found in Appendix F. We also provide some existing open-source datasets converted into GEIC format. All versions of AnythingNER are divided into train, dev, and test sets. Further details, including data volumes for different languages, are provided in Appendix B.
5 Experiment
5.1 Experiment Setting
Model English Spanish French Russian German Chinese Japanese Korean
XLM-RoBERTa 41.9 33.5 29.1 23.4 32.9 29.2 27.2 28.6
GEMNET 45.1 37.3 38.4 33.4 37.2 40.7 33.2 35.4
LS-unLLaMA 52.3 47.5 42.1 30.7 35.6 38.0 39.5 41.8
GPT-NER 46.3 40.4 38.8 34.8 39.7 35.1 36.3 36.1
CascadeNER-1.5B 58.2 52.3 51.5 49.8 47.7 54.5 52.7 51.3
CascadeNER-combine 60.2 54.2 53.4 50.5 49.1 57.4 53.2 52.5
CascadeNER-7B 60.3 55.1 54.0 51.7 50.3 58.8 54.7 52.9
Table 1: F1 score of different models across languages in the standard BIO and GEIC AnythingNER.
Model CoNLL2003 AI Literature Music Politics Science FewNERD-8 FewNERD-66
XLM-RoBERTa 92.3 59.0 65.9 72.1 70.8 66.9 80.5 64.1
ACE+document-context 94.6 17.2 22.6 23.8 35.1 32.3 83.3 70.4
BERT-MRC+DSC 93.5 63.2 67.8 74.5 76.1 68.7 86.7 74.1
LS-unLLaMA 93.1 67.1 70.2 79.2 74.2 70.4 82.0 70.9
PromptNER 84.2 64.8 74.44 84.2 78.6 72.6 76.5 35.6
GPT-NER 73.5 58.0 61.2 60.8 62.4 55.8 70.0 58.4
CascadeNER (zero-shot) 88.2 68.9 71.7 79.3 80.5 73.6 73.4 67.0
CascadeNER (few-shot) 92.8 75.8 75.2 83.2 82.4 77.1 84.5 75.9
Table 2: F1 score of different models on CoNLL2003, CrossNER, and FewNERD.
Implementation
In our experiments, we use Qwen2-1.5B and Qwen2-7B Yang et al. (2024) as the base models for CascadeNER, as they represent the current best-performing SLMs. We also evaluate Gemma, another prominent SLM, as detailed in Appendix C.2. Based on the dynamic categorized version of AnythingNER, we utilize the SWIFT framework Team (2024) to perform supervised fine-tuning (SFT) on the instruct versions of these models. Each model is fine-tuned separately on the corresponding dataset to obtain an extractor and a classifier. In addition to model sets where the extractor and classifier are of the same size, we also experiment with a configuration where the extractor uses the 1.5B model and the classifier uses the 7B model, aiming to balance accuracy and efficiency since the classification task is more complex. Due to page constraints, in experiments outside of AnythingNER, we consistently use the Qwen2-1.5B model in CascadeNER, emphasizing the framework’s low computational resource requirements and its suitability for SLMs. We evaluate CascadeNER’s performance in both few-shot and zero-shot scenarios, comparing it with supervised SOTAs and LLM-based baselines. The repetition count $i$ for result fusion is set to 3.
Dataset
We conduct experiments on the AnythingNER and existing datasets including CoNLL2003 Tjong Kim Sang and De Meulder (2003), CrossNER Liu et al. (2021), FewNERD Ding et al. (2021), PAN-X Pan et al. (2017), and MultiCoNER Malmasi et al. (2022). Since we decide to use and share these datasets in the GEIC format, we only choose open-sourced datasets to avoid copyright issues. For evaluation metrics, we primarily use F1 score, as it is widely recognized as the most robust and effective metric for NER tasks Li and Sun (2020). Potential data contamination are discussed in Appendix E.
Baselines
For supervised methods, we adopt ACE+document-context by Wang et al. (2020) (SOTA of CoNLL2003) and BERT-MRC+DSC by Li et al. (2019b) (SOTA of Ontonotes 5.0 Pradhan et al. (2013)) for English datasets, while XLM-RoBERTa Conneau et al. (2020) and GEMNET by Meng et al. (2021); Fetahu et al. (2022) (SOTA of MultiCoNER) for multilingual datasets. For LLM-based methods, we adopt GPT-NER with GPT-3.5 by Wang et al. (2023), LS-unLLaMA with LLaMA2-7B by Li et al. (2023), and PromptNER with GPT-4 by Ashok and Lipton (2023) (the current SOTA of CrossNER and FewNERD). These methods achieve the best performance when paired with the respective models according to their original papers. GPT-NER and PromptNER are few-shot learning methods, while LS-unLLaMA is a supervised method. All few-shot samples are sampled using the algorithm in Section 3.2.3. Due to the expensive price of GPT, we also applied this algorithm to limit the test set size to 300 sentences for all models. For few-shot demonstrations, PromptNER is provided with 2-shot for the best performance according to its paper, while GPT-NER and CascadeNER provided with 50-shot. Notably, when provided with more demonstrations, the performance of GPT-NER and CascadeNER improves further. We provide CascadeNER’s best performance results under the 100-shot and dual 7B model configuration in Appendix H.
Model PAN-X MultiCoNER
en es fr ru de zh ja ko en es ru de zh ko
XLM-RoBERTa 88.1 86.5 85.4 86.3 83.1 78.3 75.6 82.0 58.9 54.8 55.9 60.6 62.6 52.0
GEMNET 90.5 91.1 87.6 87.4 86.6 81.5 80.8 85.5 84.3 85.3 78.7 89.5 83.2 85.7
LS-unLLaMA 91.7 87.0 83.3 78.6 84.6 83.3 71.7 71.9 83.1 79.8 77.8 81.4 73.8 71.6
PromptNER 81.7 79.6 73.5 73.8 71.9 72.1 70.8 73.5 79.5 75.6 76.5 67.6 70.8 72.4
GPT-NER 75.2 72.8 71.6 63.5 72.0 72.4 71.5 72.1 71.7 67.9 58.2 63.1 61.2 62.5
CascadeNER (zero-shot) 87.8 85.0 83.2 80.7 77.4 78.7 74.7 72.0 71.9 71.5 71.2 63.5 70.3 69.8
CascadeNER (few-shot) 91.0 85.2 87.2 86.8 82.8 87.0 83.2 79.4 85.9 81.1 79.5 69.1 85.1 76.9
Table 3: F1 score of different models across languages on PAN-X and MultiCoNER.
5.2 Results
AnythingNER
The results are shown in Table 1. Notably, since the base models of CascadeNER are fine-tuned with AnythingNER, we only provide zero-shot results. The results indicate that, due to its unprecedentedly detailed categorization and multilingual coverage, AnythingNER is a extremely challenging flat NER dataset, placing higher demands on methods’ generalization. Additionally, its low-resource characteristics make AnythingNER more suitable for evaluating few-shot and zero-shot methods, particularly LLM-based methods. CascadeNER achieve a significant leading advantage on this dataset, demonstrating its strong generalization and multilingual proficiency. PromptNER experience serious hallucinations when faced with the complete set of 155 categories, so we fail to provide results of it.
CoNLL2003
CoNLL2003 is the most widely used English NER dataset, featuring four types: PER, LOC, ORG, and MISC. Supervised methods achieve excellent F1 scores of 90%-95% on this dataset. We use this dataset to compare CascadeNER and other LLM-based methods with existing supervised SOTAs in classical scenarios.
CrossNER
CrossNER is a English cross-domain dataset primarily used to evaluate a model’s cross-domain generalization and low-resource performance. It consists of five independent sub-datasets, each covering a specific domain (AI, Literature, Music, Politics, and Sciences) and containing 9-17 entity types. Since the train set for the datasets only contains 100-200 sentences, supervised methods underperform compared to LLM-based methods. We use this dataset to evaluate CascadeNER in cross-domain and low-resource scenarios.
FewNERD
FewNERD is an English dataset designed to evaluate a model’s ability to handle fine-grained entity recognition and few-shot learning, comprising 8 coarse-grained types and 66 fine-grained types. For supervised methods, FewNERD applies all 66 categories, challenging the models’ classification abilities. For few-shot methods, we use the Intra-10way setting, where the train, dev, and test sets contain non-overlapping entity types. We utilize both the 8-category and 66-category settings to evaluate CascadeNER under varying levels of classification granularity.
MultiCoNER & PAN-X
MultiCoNER and PAN-X are two widely used multilingual datasets. MultiCoNER covers 6 entity types and 11 languages, while PAN-X includes 3 entity types and 282 languages. We use 6 and 8 overlapping languages from MultiCoNER and PAN-X with AnythingNER to evaluate CascadeNER’s multilingual capabilities. The results are shown in Table 3. It is important to note that, for the purpose of controlling variables, all methods requiring training, including CascadeNER’s base model pre-training, are trained using multilingual joint training.
Discussion and Limitation
The results indicate that when handling NER tasks with ample training resources and simple classifications, LLM-based methods still lag behind existing methods, whether on the English-only CoNLL2003 or the multilingual PAN-X. However, in low-resource scenarios, LLM-based methods, whether supervised or few-shot, achieve significantly better results. CascadeNER surpasses existing methods on CrossNER except Music, and FewNERD, highlighting its exceptional generalization and capability to handle complex entity categorization.
However, some challenges remains. Although the GEIC task is designed to be able to accommodate nested and discontinuous NER, CascadeNER is currently limited to flat NER tasks. This limitation arises from the fact that the models in CascadeNER are pre-trained on the GEIC version of AnythingNER, and there is a significant lack of suitable multilingual corpora for nested and discontinuous NER. Our resources are insufficient to collect enough open-source data for this purpose, which lead to AnythingNER containing only flat NER labels, and thus constraining CascadeNER to flat NER. Furthermore, due to resource limitations, AnythingNER currently supports only 8 languages, excluding other cucial languages, somewhat restricting its applicability.
6 Conclusion
This paper introduces Generation-based Extraction and In-context Classification (GEIC), a novel task which is designed for applying LLMs in NER tasks, especially few-shot and zero-shot scenerios. The idea of GEIC can benefit a wide range of scenarios, not only the LLM-based NER tasks, but also inspiring researchers to apply relatively small LLMs in other complex NLP tasks, such as Relation Extraction. To demonstrate GEIC, we propose CascadeNER, the first GEIC framework, and AnythingNER, the first dataset for LLM-based NER tasks like GEIC. Extensive experiments show that CascadeNER excels in low-resource and complex classification scenarios, demonstrating strong practicality. We hope that our work will inspire further research on the application of LLMs in NLP tasks.
Appendix A Dynamic Categorization
In addition to the BIO and GEIC format with the traditional fixed categorization for AnythingNER, we introduce a GEIC version using dynamic categorization. Dynamic categorization is a novel categorization method designed for LLM-based NER training and evaluation, particularly suited for GEIC. In the dynamic categorized version of AnythingNER, we dynamically adjust the labels and corresponding entity type list for entities during annotation, overcoming the constraints of static classification. By flexibly altering the granularity and numbers of categories during data annotation, this method effectively trains and evaluates the model’s generalization. Adjustments include mixing types of different granularities, using type lists without irrelevant types, replaced types with synonyms, and and merging certain types into miscellaneous/others, as shown in Figure 6. This approach revolutionarily addresses the issue of traditional datasets being unsuitable for few-shot and zero-shot training and evaluation, simulating real-world scenarios. CascadeNER, trained on this version of the dataset, exhibits significantly improved generalization ability with extraordinary few-shot and zero-shot performance.
Figure 6: Examples of dynamic categorization methods. The red box contains the fine-grained categories of the entity in standard categorization. The green boxes contains the category lists modified with four dynamic categorization methods. The blue boxes represents the entity category corresponding to each category list.
To mitigate potential consistency issues arising from dynamic categorization, we provided annotators with detailed guidelines. Following annotation, we conduct statistical analysis and adjust the dynamically categorized entities to maintain a balance between generalization and consistency. Notably, as dynamic categorization is a subtractive approach applied to a comprehensive classification system, the success of this method relys heavily on AnythingNER’s comprehensive categorization system, which includes 155 entity types. This approach may not be suitable for all datasets.
Appendix B Detail of AnythingNER
In this section, we briefly present some key metrics of AnythingNER and compare AnythingNER with existing datasets. Table 4 presents a simple comparison between AnythingNER and existing multilingual or fine-grained datasets.
Dataset #Language #Coarse #Fine Domain
CoNLL2002 2 4 no News
CoNLL2003 2 4 no News
ACE2005 2 7 41 News
OntoNotes 5.0 3 18 no General
WNUT2017 1 6 no SocialMedia
FEW-NERD 1 8 66 General
PAN-X 282 3 no General
MultiCoNER 11 6 33 General
I2B2 1 22 no Medical
AnythingNER 8 8 155 General
Table 4: Overview of NER datasets. Notably, AnythingNER covers a wide range of cross-domain categories, such as art, medicine, and biology, thus offering better generalization compared to other general datasets.
The specific data volumes for each language are shown in Table 5. It is important to note that for languages except English and Chinese, we partially use manually translated English corpora. This is necessary to balance category distribution, as most languages lack sufficient corpora in specific domains. We also provide conversion scripts that allow AnythingNER to be transformed into train, dev, and test sets with non-overlapping subsets based on coarse categories, making it easier to use for few-shot methods.
Language # Sentences # Tokens # Entities #Train #Dev #Test
English 1500 27.3k 3709 900 300 300
Chinese 1500 26.4k 3675 900 300 300
Spanish 1000 18.6k 2393 602 201 197
French 1000 19.2k 2475 600 200 200
German 1000 18.8k 2473 603 197 200
Japanese 1000 18.4k 2387 600 199 201
Korean 1000 17.9k 2376 598 200 202
Russian 1000 19.3k 2406 601 198 201
Table 5: Statistics of AnythingNER across languages. We roughly follow a 3:1:1 ratio to divide the train, dev, and test sets, with slight adjustments based on the proportional distribution of entities within the corpus.
Appendix C Ablation Study
C.1 Result Fusion
In Section 3.2.1, we introduce our union strategy in result fusion to address the issue of extractor recall being significantly lower than precision, allowing multiple extractions for one sentence and taking the union of the results to maximize recall. For the problem of entity nesting, where different extraction rounds yield overlapping or nested entities, we adopt a length-first strategy, retaining the longer entity. Table 6 provides a example for the significantly low recall.
Dataset Precision Recall F1 Score
CoNLL2003 98.4 93.6 95.9
AI 98.7 88.0 93.1
Literature 98.3 87.8 92.7
Music 98.0 92.0 94.9
Politics 97.5 90.0 93.6
Science 98.2 85.9 91.6
Table 6: Precision, recall, and F1 Score for CoNLL2003 and CrossNER. In this experiment, both base models used in CascadeNER are Qwen2-1.5B, and the results are obtained in zero-shot scenarios.
Figure 7 presents the impact of increasing the number of extraction repetitions in zero-shot scenarios on CoNLL2003. The results show that our strategy can slightly improve recall with minimal impact on precision. Given the obvious margin effect after 3 repetitions, we ultimately select $3$ as the repetition count $k$ for other experiments. It is important to emphasize that even without repeated extraction, CascadeNER demonstrates a significant performance advantage.
Figure 7: The curves showing visualized precision, recall, and F1 Score as a function of the number of repetitions, demonstrating how these metrics change with increasing repetition counts $k$. In this experiment, both base models used in CascadeNER are Qwen2-1.5B.
C.2 Varying the Base Models
In this section, we use four different SLMs as the base models for CascadeNER, i.e., two versions of Qwen2 and Gemma with different parameters, namely Qwen2-1.5B, Qwen2-7B, Gemma-2B, and Gemma-7B. These models represent the current best-performing SLMs. We use the dynamic categorized version of AnythingNER to fine-tune the instruct versions of the four models. Each model is fine-tuned separately on the corresponding dataset to obtain both an extractor and a classifier. In addition to using combinations where the extractor and classifier are based on the same original model, we also set up combinations where the extractor uses the 1.5B/2B model and the classifier uses the 7B model, resulting in a total of six combinations. The performance comparison of these combinations on AnythingNER is shown in Table 7. Based on these results, Qwen2 outperform Gemma in multilingual tasks overall, and the performance of Qwen2-1.5B is slightly higher than Gemma-2B, while significantly reducing resource consumption. Therefore, we chose Qwen2 as the base models for other experiments.
English Chinese Spanish Japanese Macro F1
Qwen2-1.5B 71.2 67.5 65.3 65.7 67.4
Qwen2-combine 73.2 70.4 67.2 66.2 69.2
Qwen2-7B 73.3 71.8 68.1 67.7 70.2
Gemma-2B 72.9 61.5 69.1 62.4 66.5
Gemma-combine 74.6 63.3 69.7 64.2 68.0
Gemma-7B 74.8 65.9 68.3 65.1 68.5
Table 7: Comparison of F1 scores for CascadeNER using different models on AnythingNER. The results are obtained in zero-shot scenarios.
C.3 Context in Classification
In the early stages of our research, the prompt used for classification contained only the entity itself without any context. Figure 8 provides an example comparing the two types of prompts. Although this method makes the prompt more concise, it lacks any contextual information. Our final in-context classification queries significantly improve classification accuracy, as shown in Table 8.
Figure 8: Example of the early context-free queries.
Dataset ACC (context-free) ACC (in-context)
CoNLL2003 90.1 94.2
AI 75.5 79.6
Literature 78.9 83.4
Music 84.6 88.3
Politics 87.4 90.8
Science 82.2 86.5
Table 8: Both base models used in CascadeNER are Qwen2-1.5B. The results are obtained in zero-shot scenarios. The used datasets are CoNLL2003 and CrossNER. Accuracy is used in the evaluation of classifiers.
C.4 Number of Few-shot Demonstrations
In this section, we conduct experiments to estimate the effect of the number of demonstrations. We compare CascadeNER’s performance in the few-shot scenario on the CoNLL2003 dataset as the number of demonstrations increases with the current SOTA method for this dataset, ACE+document-context. The results are shown in Figure 9. While CascadeNER already demonstrate excellent performance in the zero-shot scenario, its performance improves further when provided with a certain number of few-shot demonstrations. CascadeNER shows a significant advantage in low-resource scenarios, but as the number of demonstrations increases, this advantage gradually diminishes until it is eventually surpassed by a tiny difference.
Figure 9: Comparison between CascadeNER and ACE+document-context. Both base models used in CascadeNER are Qwen2-1.5B. Demonstrations, or the sentences in the train set, are start from 0 to 10000. The sampling follows the algorithm in Section 3.2.3.
Appendix D Comparsion of Prompt
In this section, we compare our prompt with two existing LLM-based baselines, GPT-NER Wang et al. (2023) and PromptNER Ashok and Lipton (2023). These methods are the currently main methods to achieve general NER with LLMs. A breif comparsion is shown in Figure 10.
Figure 10: Examples for the three parts of the prompt for each method. The red boxes contain the task description, the green boxes contain few-shot demonstrations, and the blue boxes contain the input sentence.
PromptNER utilizes detailed descriptions of each entity’s specific definition and reasoning processes to fully leverage the LLM’s logical reasoning abilities. However, like traditional methods, it treats NER as a sequence labeling task, failing to effectively utilize the LLM’s global contextual understanding capabilities, making it prone to overlooking important context in complex sentences. Additionally, the task descriptions are overly complex, which not only makes it difficult for the model to correctly execute tasks in zero-shot scenarios, but also leads to a higher likelihood of hallucinations in tasks requiring fine-grained classification, such as the fine-grained settings of FewNERD and AnythingNER, as each category’s definition requires detailed descriptions. These issues reduce PromptNER’s generalization and accuracy, limiting its application.
GPT-NER handles the NER task by determining whether a single entity belongs to a specific category, which leverages the generative capabilities of LLMs and allows for improved attention to the influence of context on entity meaning. Its drawback lies in the fact that it can only process one entity type at a time. This makes the method highly inefficient when dealing with fine-grained categorization, leading to significant resource consumption. Additionally, this method requires multiple judgments for the same entity, introducing the potential for conflicts between different rounds. Unfortunately, GPT-NER does not provide an effective solution for this issue.
CascadeNER is based on the concept of the GEIC task, dividing the NER task into two sub-tasks: extraction and classification, while simplifying the input and output formats and reducing logical complexity. This ensures that even SLMs with limited capacity to handle complex tasks can still perform the tasks accurately and efficiently. In extraction, CascadeNER leverages the model’s generation capabilities by producing sentences with identifiers, treating all entities uniformly, which enhances the model’s generalization ability across different languages and domains. Notably, it avoids reliance on word order by consistently using "##" to mark entities, ensuring consistent annotation regardless of whether the language is right-to-left or left-to-right, improving cross-language consistency and adaptability. In classification, our method processes the entire sentence as a whole, better utilizing LLMs’ strengths in contextual understanding and semantic modeling. By leveraging the LLM’s ability to model long-range dependencies, the model’s capacity to handle complex sentence structures is enhanced, avoiding fragmentation of information and improving overall consistency and generalization. However, our method also has limitations. The use of unified identifiers prevents CascadeNER from effectively handling nested NER. We plan to address this by developing a solution that accommodates both multilingual and nested NER tasks in future.
Appendix E Data Contamination Statement
Given that LLMs are trained on data from diverse and complex sources, there is a possibility that portions of the evaluation sets may have been encountered during pre-training. However, as prior research Chowdhery et al. (2023) indicates, contaminated data that has been seen during training does not significantly influence performance. Thus, we consider this issue negligible.
Appendix F Ethical Statement of AnythingNER
When constructing AnythingNER, we strictly adhere to existing ethical guidelines Bender and Friedman (2018); Gebru et al. (2021); Hovy and Spruit (2016), ensuring that our data sources and processing methods comply with legal and ethical standards while maintaining high-quality annotations. All the text in AnythingNER is sourced from Wikipedia, ensuring no violations of privacy or copyright, as Wikipedia is an open-source platform with user-contributed content from around the world. During data collection and annotation, we balance category distribution to minimize the risk of bias in the model. Furthermore, we maintain transparency by detailing the dataset development process and data partitioning in this paper, ensuring clarity and reproducibility for future research.
For the annotators, each language in AnythingNER is annotated by two junior or higher-level students from the corresponding language departments at our university. Due to the double-blind review process, the annotators’ identities cannot be disclosed in this version. Each annotator receives extensive training and follows AnythingNER’s multi-granularity classification system to ensure consistent and accurate entity annotations across various languages and domains. The annotation process for each language are divided into two parts equally, with each annotator independently handling one part. After the initial annotation, GPT-4o OpenAI (2023) is utilized for machine review, and the annotators revise their work based on the review results. For ambiguous terms or specialized domain terms, the annotators either collaborate with each other or consult experts to ensure the accuracy and reliability of the annotations.
Appendix G Detailed Categories of AnythingNER
G.1 Person
Real Person
Politician, Artist, Author, Athlete, Director, Actor, Scholar, Military, Musician, Business Executive, Other Person.
Fictional Figure
Mythological Figure, Other Figure.
G.2 Location
Geographical Entity
Water Body, Mountain, Island, Desert, Other Geographical Entity.
Geo-Political Entity
Continent, Country, State or Province, City, District, Region, Other GPE.
Address
Address, Road, Railway, Other Address.
G.3 Product
Food
Beverages, Packaged Foods, Other Food.
Weapon
Firearms, Biological, Chemical Weapon, Explosives, Cold Weapon, Nuclear, Other Weapon.
Technology
Software, Website, Electronics, AI, Other Technology.
Vehicle
Air, Car, Water, Rail, Bike, Other Vehicle.
Other Product
Clothes, Household, Personal Care, Toys, Musical Instruments, Other Product.
G.4 Facility
Public Facility
Hospital, Library, Park, Landmark, School, Museum, Sports Facility, Other Public Facility.
Commercial Facility
Hotel, Restaurant, Market/Mall, Theater/Cinema, Bank, Other Commercial Facility.
Transportation Facility
Airport, Station, Port, Other Transportation Facility.
Production Facility
Factory, Farm, Mine, Energy, Other Production Facility.
Other Facility
Residential, Government Facility, Other Facility.
G.5 Art
Visual Art
Painting, Sculpture, Visual Art Genre, Other Visual Art.
Music
Song, Album, Music Genre, Other Music.
Literature
Poem, Non-fiction, Fiction, Literature Genre, Other Literature.
Other Art
Film, Play, Broadcast Program, Game, Other Art.
G.6 Group
Social Group
Ethnic Group, Religious Group, Other Social Group.
Non-commercial Organization
Educational and Research, Political/Military, Community, Religious Organization, Other Non-commercial Organization.
Commercial Organization
Sports Team, Band, Company, Media, Other Commercial Organization.
G.7 Miscellaneous
Award
Literary Award, Sports Award, Artistic Award, Other Award.
Event
Political/Military Event, Sporting Event, Disaster, Business Event, Other Event.
Miscellaneous
Educational Degree, Tradition, God, Law, Language, Miscellaneous.
G.8 Science Entity
Biological
Protein, Species, Biological Theory, Other Biological Entity.
Chemical
Element, Compound, Reaction, Chemical Theory, Other Chemical Entity.
Physical
Physical Phenomenon, Astronomical Object, Physical Theory, Other Physical Entity.
Computer Science
ProgramLang, Algorithm, Other Computer Science Entity.
Medical
Disease, Injury, Medication, Symptom, Medical Theory, Other Medical Entity.
Other Scientific Entity
Discipline, Academic Journal, Conference, Metrics, Other Scientific Entity.
Appendix H Best Results of CascadeNER
Model CoNLL2003 AI Literature Music Politics Science FewNERD-8 FewNERD-66
CascadeNER 93.5 76.5 76.6 83.9 83.6 78.2 86.3 76.3
Table 9: Best F1 score of CascadeNER on CoNLL2003, CrossNER, and FewNERD.
Model PAN-X MultiCoNER
en es fr ru de zh ja ko en es ru de zh ko
CascadeNER 92.1 86.1 87.5 87.1 83.8 87.9 84.4 81.4 87.1 82.2 80.7 70.5 86.2 78.3
Table 10: Best F1 score of CascadeNER on PAN-X and MultiCoNER.