### v4

# DynamicNER: A Dynamic, Multilingual, and Fine-Grained Dataset for LLM-based Named Entity Recognition

Abstract
With the advancement of Large Language Models (LLMs), more and more researchers apply LLMs for Named Entity Recognition (NER) methods, bringing vitality to this classical Natural Language Processing task. However, existing datasets are designed for traditional machine learning methods, inadequate for LLM-based methods in terms of corpus selection, entity categorization, and design logic. This limitation leads to less effective evaluation and model fine-tuning. To address this issue, we propose DynamicNER, the first NER dataset specifically designed for LLMs and with dynamic categorization, transcending the limitations of fixed categorization in existing datasets. It is also multi-lingual and multi-granular, covering 8 languages and 155 entity types, with corpus spanning multiple specialized domains. Furthermore, in response to the limitations demonstrated by existing LLM-based methods during DynamicNER testing, we develop CascadeNER, a novel NER method based on a two-stage strategy and lightweight LLMs, addressing the problems in current methods. Experiments show that DynamicNER is an effective benchmark for LLM-based NER methods, and CascadeNER outperforms existing methods with fewer computational resources. Our work is openly accessible.
DynamicNER: A Dynamic, Multilingual, and Fine-Grained Dataset for LLM-based Named Entity Recognition
Hanjun Luothanks: Zhejiang University, hanjun.21@intl.zju.edu.cn 11footnotemark: 1, Yingbin Jinthanks: Zhejiang University, 3210104412@zju.edu.cn 22footnotemark: 2, Xinfeng Lithanks: Nanyang Technological University, xinfengli@zju.edu.cn 33footnotemark: 3, Xuecheng Liuthanks: Zhejiang University, xuecheng.21@intl.zju.edu.cn 44footnotemark: 4, Ruizhe Chenthanks: Zhejiang University, ruizhec.21@intl.zju.edu.cn 55footnotemark: 5, Tong Shangthanks: Southeast University, 213211756@seu.edu.cn 66footnotemark: 6, Kun Wangthanks: Nanyang Technological University, wk520529@mail.ustc.edu.cn 77footnotemark: 7, Qingsong Wenthanks: Squirrel AI, qingsongedu@gmail.com 88footnotemark: 8, Zuozhu Liuthanks: Zhejiang University, zuozhuliu@intl.zju.edu.cn, Corresponding author 99footnotemark: 9
1 Introduction
The rapid advancement of Large Language Models (LLMs) brings revolutions to natural language processing (NLP) (Naveed et al., 2023). As a classical NLP task, Named Entity Recognition (NER) also witnesses significant method innovations driven by LLMs Xie et al. (2023). Existing LLM-based methods Shao et al. (2023); Li and Zhang (2023), leveraging LLMs’ generalization and contextual understanding capabilities, show superior performance compared to traditional machine learning (ML) methods in low-resource, multilingual, or few/zero-shot learning scenarios. This advantage is particularly crucial in AI for Science and Healthcare, where domain resources, especially non-English materials, are often limited. Thus, LLM-based methods gain attention from researchers in these fields Xiao et al. (2024).
Nevertheless, there are no existing NER datasets optimized for the characteristics of LLMs, limiting the effective training and evaluation of LLM-based methods. This limitation manifests primarily in three aspects. First, existing datasets employ static categorization with a fixed set of entity types, preventing effective evaluation of LLMs’ generalization capabilities in handling new entity types and varying granularities under few-shot/zero-shot scenarios. Secondly, the corpus in existing NER datasets focuses on short texts and isolated sentences, making them ineffective to evaluate the capabilities of LLMs in long-range context. Finally, while existing datasets separately address specialized corpus with corresponding entity types Kim et al. (2003); Liu et al. (2021), multi-grained classifications Ding et al. (2021), and multi-languages Malmasi et al. (2022), no dataset simultaneously incorporates three aspects, limiting comprehensive evaluation of LLM-based methods, which show potential across all these aspects. As a result, existing datasets lack the capability to effectively discriminate performance differences among LLM-based methods, fail to reflect their true potential and limitations, and impedes future improvements.
To address this issue, we develop DynamicNER, the first NER dataset optimized for LLM-based methods and the first NER dataset with dynamic categorization. It employs multiple strategies to dynamically adjust entity labels, type lists, and granularity levels during annotation, enhancing the evaluation of generalization ability of NER methods. We design cohesion and distribution balance metrics to enable evaluation and optimization during the categorization, with the entire procedure being algorithmically automated, ensuring reliability and reproducibility. This method addresses the limitations of existing datasets in training and evaluating few-shot/zero-shot learning ability. Besides, DynamicNER is a multilingual and multi-granular dataset, featuring 8 languages, 8 coarse-grained types, 31 medium-grained types, and 155 fine-grained types, with entity types and corpora covering extensive professional content across domains including sciences, medicine, and art, providing an unprecedented level of coverage.
Furthermore, our test on DynamicNER reveals significant limitations in existing methods, particularly when migrating to lightweight LLMs (LLMs with parameters from 1.5B to 7B) for local deployment. While methods with commercial models like GPT Brown (2020) achieve reported performance levels, this dependency introduces challenges regarding API costs and privacy risks. Using models via API are expensive for practical NER application. Privacy concern is also a significant issue Zhang et al. (2024); Das et al. (2024); Deng et al. (2025). Given the current absence of laws regarding LLM data uploading in many countries, users face difficulties in protecting their rights. Therefore, preventive measures to avoid unauthorized disclosure of information becomes essential.
For this issue, we propose CascadeNER, a universal and multilingual NER framework that achieves comparable performance with lightweight LLMs to existing LLM-based methods requiring expensive commercial models. CascadeNER employs a two-stage strategy, redefines NER as two in-context text generation sub-tasks, extraction and classification. By model cascading Varshney and Baral (2022), CascadeNER assigns each sub-tasks to corresponding fine-tuned lightweight LLMs, decreasing task complexity and helping capture in-context dependencies. Leveraging the prior knowledge, CascadeNER can complete NER tasks in multiple languages in low-resource scenarios effectively.
We evaluate a BERT-based Devlin et al. (2018) supervised method, two LLM-based methods, and our proposed CascadeNER on DynamicNER. The three LLM-based methods are tested on 3 difference models: Qwen2.5-1.5B, Qwen2.5-7B Bai et al. (2023), and GPT-4o Achiam et al. (2023). We also conduct evaluations of CascadeNER against existing methods on existing datasets. Results demonstrate that DynamicNER effectively evaluates the performance of LLM-based methods in low-resource and complex NER tasks, while CascadeNER outperforms existing LLM-based methods significantly with smaller models. Moreover, this work offers the first comprehensive comparison and analysis of existing LLM-based NER methods, with a emphasis on multilingual and fine-grained scenarios.
Our core contributions are summarized as follows:
  • We build DynamicNER, the first NER dataset optimized for LLM-based NER with a novel dynamic categorization system, which supports 8 languages, 155 entity types, and three-level granularity
  • We propose CascadeNER, a universal and multilingual NER framework, which outperforms existing LLM-based methods with lightweight LLMs
  • We conduct the first comprehensive comparison of LLM-based NER methods, discussing future directions of LLM-based NER
2 Related Works
Named Entity Recognition
NER is the task of identifying named entities in text and classifying them into predefined categories. Supervised methods, such as BiLSTM Yu et al. (2020) and BERT-MRC Li et al. (2019a), currently dominate this task. They generally rely on large amounts of train data to achieve strong performance, which limits their application in low-resource scenarios. Some researchers apply LLMs to address this issue. GPT-NER Wang et al. (2023) employs the GPT-3 model and re-frames the task as single-entity labeling, supporting few-shot/zero-shot learning. It achieves comparable performance to supervised methods in traditional scenarios and excels in low-resource scenarios. PromptNER Ashok and Lipton (2023) achieves state-of-the-art (SOTA) accuracy in datasets with complex classification Liu et al. (2021); Ding et al. (2021) with GPT-4 and Chain-of-Thought (CoT) Wei et al. (2022), yet performs significantly worse than GPT-NER and supervised methods in classical NER datasets like CoNLL2003 Tjong Kim Sang and De Meulder (2003). Furthermore, several studies apply LLM-based NER in domain-specific tasks Li and Zhang (2023); Shao et al. (2023); Keloth et al. (2024), focusing on science and medicine. Their performances surpass supervised methods in those domains, further highlighting the potential of LLMs in low-resource and complex NER tasks.
NER Dataset
There have been a considerable number of NER datasets in various domains Tjong Kim Sang (2002); Kim et al. (2003); Doddington et al. (2004); Walker et al. (2006); Weischedel et al. (2011); Pradhan et al. (2013); Derczynski et al. (2017). However, these existing datasets exhibit several limitations, making them unsuitable for LLM-based NER. Most previous multilingual NER datasets adopt coarse-grained classification, no longer meeting the fine-grained requirements of contemporary flat NER applications. Even existing fine-grained datasets demonstrate clear limitations in category coverage and granularity, falling short of being truly "universal." For instance, FewNERD, despite having 66 entity types, suffers from highly imbalanced data distribution, which affects its reliability for evaluating few-shot learning capabilities. Furthermore, current datasets fail to adequately address the generalization capabilities of LLMs, hindering the comprehensive training and evaluation of LLM-based NER methods. Table 1 presents a simple comparison between DynamicNER and existing multilingual or fine-grained datasets.
Dataset #Language #Coarse #Fine Domain
CoNLL2002 2 4 no News
CoNLL2003 2 4 no News
ACE2005 2 7 41 News
OntoNotes 5.0 3 18 no General
CrossNER 1 9-17 no Multi Domain
FEW-NERD 1 8 66 General
PAN-X 282 3 no General
MultiCoNER 11 6 33 General
I2B2 1 22 no Medical
DynamicNER 8 8 155 Multi Domain
Table 1: Overview of NER datasets. Notably, DynamicNER covers a wide range of cross-domain categories, such as art, medicine, and biology, thus offering better generalization compared to other general datasets.
3 DynamicNER Dataset
DynamicNER spans 8 languages: English, Chinese, Spanish, French, German, Japanese, Korean, and Russian. In terms of categorization, it is the first NER dataset with three-level granularity categorization, encompassing 8 coarse-grained types, 31 medium-grained types, and 155 fine-grained types, as shown in Figure 1. Like other NER datasets, DynamicNER is divided into train, dev, and test sets, and data volumes for different languages and parts shown in Appendix C. To develop DynamicNER, we first collect unlabeled corpus from Wikipedia. Then we manually extract sentences from corpora and annotate entities. After human annotation, we guide the dynamic categorization with category cohesion and distribution uniformity to automatically process base DynamicNER, and results in one base version and one dynamic version. Details are given in the following parts.
Figure 1: The coarse-grained and medium-grained categories of DynamicNER. Detailed categories are provided in Appendix J.
Corpora Collection and Annotation
Wikipedia provides multilingual, domain-specific corpora with clear hierarchical and indexing systems, serving as a rich resource for our research. We utilize legal Wikipedia-API to filter and download corpora across different languages and categories, followed by manual selection and annotation of sentences. We particularly focus on corpora containing long texts and complex contexts. After completing 50% of the annotation process of each language, we annotate corpora from categories related to underrepresented entity types to achieve a balanced entity type distribution. For instance, when the entities of "Algorithm" are significantly less than others, we use more corpora from Computer Science category. As a result, DynamicNER ensures balanced entity distribution, and includes rare entities and emerging fields which are not adequate in existing datasets, ensuring comprehensive coverage across diverse domains. A statement about annotators and ethical considerations is provided in Appendix I.
Dynamic Categorization
The dynamic version improves model generalization and reduces overfitting risk by dynamically adjusting entity labels and corresponding entity type lists during annotation, including mixing types of different granularities, replacing types with synonyms, using type lists without irrelevant types, and merging certain types into miscellaneous/others, as shown in Figure 2. This method addresses the mismatch between existing datasets and few-shot/zero-shot training needs, better simulating real-world scenarios, and is particularly critical for evaluating methods relying on complex prompt designs (e.g., CoT). Unlike traditional few-shot learning, some LLM-based methods only use few-shot demos to help the model understand the task or format, without requiring knowledge of entity types. They can perform NER across different datasets with fixed few-shot demos, resembling zero-shot NER. Research shows this method is more effective than typical zero-shot NER Zhang et al. (2022). In methods that uses complex prompt designs like CoT to guide the reasoning, even few-shot CoT only conveys the CoT process rather than task-relevant knowledge, the performance of prompt-guide zero-shot CoT is significantly worse than few-shot CoT, making zero-shot restrictions inadequate for reflecting their true capability. However, in NER, models inevitably learn about entity types through few-shot demos, which limits generalization evaluations on fixed-category datasets. Our method significantly mitigates this limitation by varying entity types and lists, isolating the impact of prior type knowledge.
Figure 2: Examples of dynamic categorization. The red box contains the fine-grained categories in base version. Green boxes contains the lists modified with 4 methods. Blue boxes contains the entity type to each list.
Notably, as dynamic categorization is a subtractive method applied to a comprehensive classification system, this method highly relies on DynamicNER’s comprehensive categorization system, which includes 155 entity types. This method may not be suitable for all datasets.
Categorization Metrics
Random dynamic categorization not only exhibits poor reproducibility and explainability, but may also lead to data quality degradation. For training, inappropriate categorization may result in inconsistent learning objectives and overfitting risks Ren et al. (2016). For evaluation, certain categories may experience imbalanced sample distribution and boundary ambiguity, reducing the comprehensiveness and consistency of evaluation Obeidat et al. (2019). Thus, we design four metrics to regulate the dynamic categorization: cohesion, normalized entropy, Gini coefficient, and variation coefficient. The definition and calculation methods are provided in Appendix A.
Figure 3: Pipeline of dynamic categorization. Blue boxes contains methods of each re-categorization. Yellow boxes contains used metrics of each evaluation.
Categorization Process
The dynamic categorization process consists of 4 rounds of re-categorization, each sequentially corresponding to an adjustment method, and different metrics are employed in each round to guide the optimization. This hierarchical design enables each stage to focus on distinct data characteristics and optimization objectives, preventing interference between metrics while ensuring proper optimization direction, thus achieving a progressive optimization. We do not use all metrics in each evaluation, considering that certain metrics may have overlapping or conflicting effects at specific stages. For instance, normalized entropy and Gini coefficient both measure distribution uniformity, while improving cohesion may lead to more concentrated distribution and consequently lower entropy values. Figure 3 illustrates the metrics and methods corresponding to each round. Appendix B explains the reasons of metric selection for each round.
4 CascadeNER
Figure 4: Use a sentence and the multi-granularity categories of DynamicNER as the example. The extractor and classifier are the two different lightweight LLMs used in CascadeNER. Azure boxes represent the specific type list for the input of the classifier. Blue boxes represent the sentence input.
4.1 Framework
Background
Some existing supervised methods suggest that separating extraction and classification can improve NER performance as this two-stage strategy reduces the task complexity Shen et al. (2021); Wu et al. (2022). However, these methods are limited by traditional models, failing to incorporate LLMs, and exhibit notable performance deficiencies that make them inferior to other methods treating NER as a single task. On the other hand, LLM-based methods demonstrate superior performance compared to traditional methods in Named Entity Extraction Sancheti et al. (2024) and Text Classification Gasparetto et al. (2022), indicating the potential of two-stage in LLM-based NER.
Framework Design
We propose the framework to implement two-stage strategy in LLM-based NER. CascadeNER divides NER into two sequential, independently executed, generation-based sub-tasks. In the first sub-task, extraction, the model generate a sentence where all named entities are marked with identifiers and individually re-embeds each entity back into its context, resulting in sentences with identifiers at the number of entities. In the second sub-task, classification, the model receives sentences with identifiers and a list of entity types, and label one entity at a time.
Model Cascading
To optimize performance while reducing computational resources, CascadeNER employs model cascading, where the extraction and classification sub-tasks are handled separately by two specialized fine-tuned LLMs. This structure allows each model to focus on its specific sub-task, maximizing performance on simpler, more specialized tasks. The architecture enables CascadeNER to be particularly suitable for lightweight LLMs, as each model only focus on a simplified task. Existing research shows that fine-tuned lightweight LLMs can achieve performance close to normal LLMs on specific simple tasks Hu et al. (2024a). Through the implementation of two-stage strategy and model cascading, CascadeNER effectively leverages the advantages of lightweight LLMs in simple tasks, maintaining high accuracy while reducing computational resource usage.
Pipeline
A simplified pipeline of CascadeNER with an example is shown in Figure 4. Upon receiving the input sentence, CascadeNER first processes the sentence by the extractor to mark all entities with identifiers, and re-embeds each entity back, resulting in sentences with identifiers around the named entities. These sentences are then individually fed into the classifier, which classifies each entity based on the context and the input type list. For multi-granularity data, CascadeNER allows a progressive strategy, significantly improving CasacdeNER’s performance in accurate fine-grained classification. The detailed steps of extraction and classification are discussed in following sections.
4.2 Extraction
Prompt Design
In the extraction sub-task, we utilize a generation-based extraction method, where special tokens "##" are used to surround any entities identified in the sentence, regardless of the number of entities or their types. For example:
Q: Apple proposes new Macbook
A: ##Apple## proposes new ##Macbook##
This method, compared to conventional sequential extraction, avoids requiring LLMs to perform text alignment, thus reducing task complexity. Comparing similar methods Wang et al. (2023); Hu et al. (2024b), CascadeNER’s query contains only the sentence, without any task descriptions, demonstrations, or category information. The response exclusively uses "##" as the identifiers, and all entities are extracted without specifying categories. CascadeNER achieves low-cost NER by using simple prompts and better generalization by treating all entities uniformly. A detailed comparison with existing methods and further advantages of our method are shown in Appendix F.
Result Fusion
After conducting extensive experiments, we find that the extractor’s precision consistently exceeds recal, regardless of the model or dataset, indicating that while correct entities are effectively identified, there is a tendency for under-detection. To mitigate this issue, we introduce a union strategy in result fusion Ganaie et al. (2022), allowing multiple extraction for one sentence and taking the union of the results to maximize recall. For cases of entity nesting, where different extraction rounds produce overlapping or nested entities, we apply a length-first strategy, retaining the longer entity, as longer entities generally carry more semantic meaning Nguyen and Cao (2010). For example, "Boston University" is semantically more accurate than "Boston" in the context of "She studies in Boston University". The formula of our strategy is shown below:
$E_{\text{final}}=\bigcup_{i=1}^{n}\left\{\arg\max_{e\in E_{i}}\text{length}(e)\right\}$ (1)
where $E_{i}$ is the set of extracted entities from the $i$-th extraction, $n$ is the number of extraction rounds, $E_{\text{overlap}}$ is the set of overlapping or nested entities, and $\text{length}(e)$ is the length of entity $e$. The effects of the number of extraction repetitions and other details are provided in Appendix E.1.
4.3 Classification
Prompt Design
In the classification sub-task, we employ a generation-based in-context classification method, where we input the categories and the sentence with one entity surrounded by "##", and require the classifier to generate the label for that entity. This method re-embeds the entity into the sentence for classification, which utilizes the self-attention architecture of LLMs for contextual information and improves accuracy compared to entity-level classification. Figure 5 is an example:
Figure 5: Example prompts of classification.
In zero-shot scenarios, we use a slightly different prompt. Due to differences in entity categorization across datasets, some entities in one dataset may be overlooked in others. We append the query with If none of them applied, return unknown to handle situations where the extracted entity cannot be classified into the provided categories, enhancing CascadeNER’s generalization.
Multi-granularity
For multi-granularity data, we apply a progressive strategy. After obtaining the coarse-grained result, CascadeNER use the result to index the corresponding sub-categories and classify again, continuing this process until no further classification is possible:
$L_{i}^{\text{fine}}=f_{\text{fine-classify}}(L_{i}^{\text{coarse}},\text{subcategories})$ (2)
where $L_{i}^{\text{fine}}$ is the fine-grained label, $L_{i}^{\text{coarse}}$ is the generated coarse-grained label, and subcategories are the subcategories under the coarse-grained.
5 Experiment
In this section, we first present the categorization metric changes of DynamicNER before and after dynamic categorization, followed by a comparative analysis of existing methods and CascadeNER’s performance on different versions of DynamicNER. We also conduct experiments of CascadeNER and baselines on existing datasets, along with ablation studies. In evaluations across existing datasets, CascadeNER, with base models fine-tuned using the dynamic version of DynamicNER, demonstrates consistent excellence in all datasets and achieves new SOTA performance in both FewNERD and CrossNER datasets. This confirms that DynamicNER not only provides exceptional effectiveness for evaluating LLM-based NER methods but also offers substantial value in training. Due to space limitations, these experiments are provided in Appendix D and E.
Figure 6: Quantitative categorization metric results for 3 versions DynamicNER in English and Chinese. Generally, higher cohesion and normalized entropy, or lower Gini coefficient and variation coefficient, indicate better quality.
5.1 Categorization Quality Evaluation
To demonstrate that our dynamic categorization improves dataset generalization while maintaining dataset quality, we conduct comparative experiments across three versions of DynamicNER: the Base Version, a version with random parameters for dynamic categorization, and the optimized Dynamic Version. We still employ the 4 metrics for evaluating dataset quality, whose detailed definition are provided in Appendix A. For reliability of the random version’s results, we conduct five independent tests and use the average results. Due to space limitations, we only present results for English and Chinese in Figure 6 here. Other quantitative results are provided in Appendix K.
Experimental results demonstrate that our dynamic categorization significantly increases data diversity, as shown in Table 2, while maintaining or improving dataset quality compared to the base version. The quality of the dynamic version also considerably surpasses the random version. These results comprehensively validate the reliability and effectiveness of our method.
Language en es fr ru de zh ja kr
# Lists 725 455 501 377 465 786 553 478
Table 2: The numbers of entity type lists of each language after dynamic categorization. In some scenarios, this can be equivalent to having 700+ distinct datasets.
5.2 DynamicNER Experiment
Baseline Selection
In our experiments for DynamicNER, we evaluate four NER methods: XLM-RoBERTa Conneau et al. (2020), GPT-NER Wang et al. (2023), PromptNER Ashok and Lipton (2023), and our CascadeNER. XLM-RoBERTa is a famous BERT-based multilingual model widely used as a baseline in multilingual NER research Malmasi et al. (2022); Fetahu et al. (2022), thus being selected as our baseline representing supervised methods. GPT-NER and PromptNER are two major general LLM-based NER methods that achieve performance significantly superior to supervised methods in low-resource scenarios through sophisticated prompt design and powerful GPT models, as discussed in Section 2 and Appendix F.
Model Selection
Given the lack of existing lightweight LLM-based NER methods and controlled variable principles, we evaluate two LLM-based methods and CascadeNER using three LLMs: Qwen2.5-1.5B, Qwen2.5-7B, and GPT-4o. Qwen series is the largest open-source model family. Its lightweight LLMs perform exceptionally across benchmarks and gaining widespread recognition. According to HuggingFace (2024), Qwen2.5-1.5B is the most downloaded open-source model in 2024. Therefore, we select Qwen2.5-1.5B and 7B to represent the current best-performing lightweight LLMs. GPT-4o is the most widely-used current general commercial LLM, and its previous versions are employed in GPT-NER and PromptNER, making it our representative choice for commercial LLMs. In CascadeNER, for statistics convenience, the extractor and the classifier use the same base model.
Model Dynamic-Supervised Dynamic-Fewshot
en es fr ru de zh ja ko en es fr ru de zh ja ko
G-1.5B 47.6 39.7 38.0 37.6 37.3 41.2 35.7 36.1 36.9 32.2 31.9 30.5 30.3 35.8 31.9 32.6
G-7B 52.3 46.4 44.8 44.8 45.7 48.1 42.3 42.1 42.7 37.3 38.2 36.8 36.5 41.1 37.3 38.6
G-GPT 60.6 57.3 56.5 55.6 55.9 58.4 54.9 53.8 49.2 46.9 47.5 47.2 47.0 48.9 47.7 48.3
P-1.5B 23.2 20.8 18.5 16.3 17.5 22.7 18.0 17.3 20.5 17.9 16.2 15.9 16.1 19.9 16.0 15.9
P-7B 44.3 35.8 33.2 32.5 31.9 40.4 37.4 35.6 39.8 33.2 32.1 31.8 31.5 37.8 35.6 34.5
P-GPT 53.0 50.5 51.2 47.9 50.2 52.3 48.7 48.5 49.4 48.5 47.1 46.6 46.0 47.4 44.1 44.0
C-1.5B 62.8 55.7 52.8 51.1 48.8 58.9 54.1 52.7 49.7 44.1 44.0 43.4 42.9 48.5 43.1 43.8
C-7B 68.2 61.5 55.3 52.9 51.4 64.5 58.8 55.3 55.7 49.9 49.7 46.5 46.1 52.9 50.2 50.0
C-GPT 73.1 67.1 67.8 66.9 67.6 68.3 67.4 67.9 61.3 57.4 56.9 56.2 56.0 59.7 56.8 56.4
Table 3: The results of supervised learning with dynamic version and few-shot learning with dynamic version. G means GPT-NER, P means PromptNER, and C means CascadeNER. The results indicate that, due to its unprecedentedly detailed categorization and multilingual coverage, DynamicNER is a extremely challenging flat NER dataset, placing higher demands on methods’ generalization capability.
Implementation
For the supervised method, XLM-RoBERTa is trained and only evaluated with the base version of DynamicNER. As its fixed classification output layer corresponds to a predefined set of entity types and any modification to the entity type list necessitates full model retraining, it can not be evaluated with the dynamic version. For LLM-based methods, we conduct experiments under three scenarios: supervised learning with base version, supervised learning with dynamic version, and few-shot learning with dynamic version. Training data for GPT-NER and CascadeNER is obtained through format conversion. For PromptNER, as its prompt involves complex designs such as CoT, we utilize LLM-generated prompts by GPT-4o using prompts from its paper as few-shot demonstrations and manually verified the prompts. The repetition count $i$ of CascadeNER for result fusion is set to 3. Potential data contamination are discussed in Appendix H.
Model Base-Supervised
en es fr ru de zh ja ko
BERT 41.9 33.5 29.1 23.4 32.9 29.2 27.2 28.6
G-1.5B 50.2 43.5 40.4 39.8 39.3 44.1 38.9 38.7
G-7B 55.1 48.2 47.2 44.0 48.1 50.9 44.8 44.5
G-GPT 62.4 58.3 57.9 56.8 56.9 60.4 57.3 55.9
P-1.5B 21.6 18.6 17.1 14.9 15.8 20.7 16.4 15.9
P-7B 41.1 32.9 31.0 30.7 30.3 47.4 35.6 29.6
P-GPT 49.7 47.7 48.2 45.9 46.6 48.6 45.7 45.4
C-1.5B 67.6 59.9 57.9 55.7 53.5 64.0 58.5 55.1
C-7B 73.8 65.5 60.3 59.6 61.4 69.8 65.3 62.7
C-GPT 77.1 71.7 69.9 70.3 70.8 74.3 72.4 70.9
Table 4: The results of supervised learning with base version. RoBERT represents XLM-RoBERTa. The supervised method XLM-RoBERTa performs terribly.
Results and Discussion
The results are shown in Table 3 and 4. CascadeNER achieve a significant advantage on DynamicNER, demonstrating its strong generalization and multilingual proficiency. The supervised method XLM-RoBERTa performs terribly, as DynamicNER’s low-resource characteristics make it more suitable for evaluating LLM-based methods. For LLM-based methods, the 3 methods show significant performance variations across different datasets and models. When using GPT-4o and transitioning from supervised to few-shot, PromptNER exhibits notably smaller performance degradation, partially reflecting the generalization advantages of reasoning-focused approaches. However, when migrating to lightweight LLMs, these methods show significantly larger performance drops compared to the other two methods. GPT-NER and CascadeNER demonstrate generally similar performance patterns, but GPT-NER shows more pronounced degradation when migrated to lightweight LLMs, while CascadeNER achieves a greater performance advantage on the dynamic version compared to the base version, validating the effectiveness of the two-stage strategy in both complex classification and reducing computional resource requirement.
6 Conclusion
This paper introduces DynamicNER, a multi-lingual and multi-granular NER dataset optimized for LLM-based NER, including a human-annotated base version and a dynamic-categorized version. We develop the first dynamic categorization method in NER datasets for DynamicNER, enhancing its generalization while keeping data quality. We also propose CascadeNER, a powerful NER framework which is exceptionally suitable for lightweight LLMs and local deployment, outperforming current LLM-based methods. Moreover, we conduct comprehensive experiments and analyses on DynamicNER and discuss the advantage and future direction of LLM-based NER. More experiments and discussions are provided in Appendix D and F. We hope that DynamicNER and CascadeNER will facilitate future research in LLM-based NER, revitalizing this classical NLP task.
7 Limitation
There are still some challenges for our research. Although CascadeNER is designed to be able to accommodate nested and discontinuous NER, we only conduct evaluation on CascadeNER about flat NER tasks. This limitation arises from the fact that the models in CascadeNER are pre-trained on the dynamic version of DynamicNER, and DynamicNER is a flat NER dataset. Our resources are insufficient to collect enough open-source data for this purpose, which lead to DynamicNER containing only flat NER labels, and thus constraining CascadeNER to flat NER. Furthermore, Due to resource constraints and our failure to find annotators proficient in other languages for manual annotation, DynamicNER currently supports only 8 languages, which somewhat restricts its applicability.
Appendix A Categorization Metric Definition
Cohesion
Category cohesion score (cohesion) measures categorical semantic consistency by calculating the average semantic similarity between all entities within the same category. We employ the BERT-base Devlin et al. (2018) model to extract semantic representations of entities, obtain embeddings of each entity, then computing cosine similarity between embeddings to derive cohesion. This metric ranges from [-1,1], where 1 indicates complete similarity and -1 indicates complete opposition. Typically, we perform category merging when cohesion exceeds 0.9. The formula is shown below:
$Cohesion=\frac{1}{n(n-1)}\sum_{i=1}^{n}\sum_{j=i+1}^{n}\cos(\mathbf{v}_{i},\mathbf{v}_{j})$ (3)
where $n$ is the number of entities in this category, $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ are the vector representations of the $i$-th and $j$-th entities encoded by BERT-base, and $\cos(\mathbf{v}_{i},\mathbf{v}_{j})$ represents the cosine similarity between two vectors.
The detailed cosine similarity formula is shown below.
$\cos(\mathbf{v}_{i},\mathbf{v}_{j})=\frac{\mathbf{v}_{i}\cdot\mathbf{v}_{j}}{\|\mathbf{v}_{i}\|\|\mathbf{v}_{j}\|}$ (4)
Normalized Entropy
Normalized entropy measures the overall balance of category distribution. This metric is used for the influence of category quantity by calculating the information entropy of category frequency distribution and normalizing it to a score within the range [0,1]. A score of 1 indicates perfect balance, where all categories have equal sample sizes, while 0 indicates complete imbalance, where all samples are concentrated in a single category. When normalized entropy falls below 0.8, it indicates significant distributional imbalance and needs to be adjusted. The formula is shown below:
$H=-\frac{\sum_{i=1}^{n}p_{i}\log_{2}(p_{i})}{\log_{2}(n)}$ (5)
where $n$ is the total number of categories, $p_{i}$ is the proportion of samples in the $i$-th category calculated as the number of samples in category $i$ divided by the total number of samples across all categories.
Gini Coefficient
The Gini Coefficient Gini (1921) measures the degree of inequality in category distribution. Compared to normalized entropy, the Gini coefficient demonstrates higher sensitivity to distributional inequalities, performing better at identifying extreme imbalances where minority categories contain large sample proportions. For instance, when sample distributions exhibit extreme imbalances like [0.8, 0.1, 0.05, 0.05], the Gini coefficient provides stronger warning signals, while normalized entropy is more suitable for monitoring progressive imbalances such as [0.4, 0.3, 0.2, 0.1]. This metric also ranges from [0,1], where 0 indicates perfect balance and 1 indicates complete imbalance. A Gini coefficient exceeding 0.4 signals significant categorical inequality and requires distribution improvement. By using Gini coefficient and normalized entropy together, we achieve both sensitive detection of extreme imbalances and effective monitoring of overall distribution trends. The formula is shown below:
$G=\frac{n+1-2\sum_{i=1}^{n}(n-i+1)p_{i}}{n}$ (6)
where $n$ is the total number of categories, $p_{i}$ is the proportion of samples in the $i$-th category after sorting proportions in ascending order $(p_{1}\leq p_{2}\leq...\leq p_{n})$.
Variation Coefficient
The Coefficient of Variation measures data dispersion by calculating the ratio of standard deviation to mean of category sample sizes. Its advantage is its scale independence, enabling comparisons across different scenarios. The coefficient ranges from 0 to positive infinity, where 0 indicates perfect balance and larger values indicate greater distributional imbalance. When the coefficient exceeds 0.5, it indicates significant fluctuation in sample sizes between categories, necessitating balance adjustments. The formula is shown below:
$CV=\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}}{\bar{x}}=\frac{\sigma}{\mu}$ (7)
where $n$ is the total number of categories, $x_{i}$ is the number of samples in the $i$-th category, $\bar{x}$ is the mean number of samples across categories, $\sigma$ is the standard deviation of sample numbers, and $\mu$ is the mean.
Appendix B Categorization Metric Selection
Mixing Types of Different Granularities
In this round of re-categorization, we use cohesion, normalized entropy, and Gini coefficient as metrics for optimization. Cohesion is employed to assess relationships between entity types, where close categorical relationships reduce the need for mixing to avoid creating unreasonable combinations. Meanwhile, normalized entropy and Gini coefficient are utilized to comprehensively measure distribution uniformity, where uneven distributions guide the system to perform additional merging for balance or category redistribution.
Replace with Synonyms
In this round of re-categorization, we use Gini coefficient and variation coefficient as metrics for optimization. We employ the variation coefficient to measure data dispersion, increasing synonym substitutions for increasing data convergence when dispersion is high. The Gini coefficient is used to guide system to reduce operations to prevent exacerbating imbalances when distributions are uneven. Cohesion is not used as synonym substitution does not alter hierarchical relationships between categories. Entropy is also given up because synonym substitution primarily focuses on linguistic variation rather than distributional changes.
Remove Irrelevant Types
In this round of re-categorization, we use cohesion and normalized entropy as metrics for optimization. We employ normalized entropy as a reference for controlling removal probability, ensuring that deletion operations do not result in overly concentrated distributions. Additionally, the system adjusts removal probability when cohesion is low, regulating relationships between categories. The variation coefficient is not used as this stage primarily focuses on option quantity rather than distribution characteristics, while the Gini coefficient is omitted since distribution balance has been addressed in the previous two stages, thus temporarily foregoing the Gini coefficient to prevent interference with other metrics.
Merge Types into Miscellaneous
In this round of re-categorization, we use all four metrics for optimization. As the final optimization stage, it requires consideration across all dimensions. We use all metrics for final fine-tuning to ensure overall data quality and avoid biases that might arise from single metric optimization.
Appendix C Detail of DynamicNER
The specific data volumes for each language are shown in Table 5. It is important to note that for languages except English and Chinese, we partially use manually translated English corpora. This is necessary to balance category distribution, as some languages lack sufficient corpora in specific domains. We also provide conversion scripts that allow DynamicNER to be transformed into train, dev, and test sets with non-overlapping subsets based on coarse categories, making it easier to use for traditional few-shot learning methods.
Additionally, two points about the Table 5 require clarification. First, as DynamicNER’s design emphasizes the evaluation of generalization and low-resource learning capabilities, we set the test set capacity to the biggest one, rather than the train set. Second, for Chinese, Japanese, and Korean, due to linguistic characteristics where each character is treated as a token, the token count appears significantly higher, though the actual corpus volume is comparable to other languages.
Language # Sentences # Tokens # Entities # Train # Dev # Test
English 1500 36.7k 4664 300 300 900
Chinese 1500 98.1k 5198 300 300 900
Spanish 1000 22.8k 2454 197 201 602
French 1000 24.1k 2763 200 200 600
German 1000 21.7k 2800 200 197 603
Japanese 1000 81.7k 3032 201 199 600
Korean 1000 66.4k 2401 202 200 598
Russian 1000 18.5k 2092 201 198 601
Table 5: Statistics of DynamicNER across languages. We roughly follow a 1:1:3 ratio to divide the train, dev, and test sets, with slight adjustments based on the proportional distribution of entities within the corpus.
Appendix D More Experiment about CascadeNER
D.1 CascadeNER Setting
In the experiments of this section, CascadeNER always employs two Qwen2.5-7B base models, which are fine-tuned separately based on the corresponding part of the dynamic version of DynamicNER to obtain an extractor and a classifier. Potential data contamination about the fine-tuning is discussed in Appendix H. We evaluate CascadeNER’s performance in both few-shot and zero-shot scenarios, comparing it with supervised SOTAs and LLM-based baselines. For few-shot scenarios, the number of few-shot demonstrations is set to $3$, the same as the experiments on DynamicNER.
D.2 Baselines
For supervised methods, we adopt ACE+document-context by Wang et al. (2020) (SOTA of CoNLL2003) and BERT-MRC+DSC by Li et al. (2019b) (SOTA of Ontonotes 5.0 Pradhan et al. (2013)) for English datasets, while XLM-RoBERTa Conneau et al. (2020) and GEMNET by Meng et al. (2021); Fetahu et al. (2022) (SOTA of MultiCoNER) for multilingual datasets. For LLM-based methods, we adopt GPT-NER and PromptNER with GPT-4o.
D.3 Dataset
Few-shot Data Sampling
In existing datasets, only CrossNER Liu et al. (2021), designed for low-resource scenarios, and FewNERD Ding et al. (2021), designed for few-shot scenarios, meet our requirements for evaluating CascadeNER in few-shot scenarios. However, relying solely on them is insufficient for comprehensively evaluating CascadeNER, particularly its multilingual NER performance. To address this, we develop a sampling algorithm to construct datasets for few-shot evaluation. Considering that basic random sampling cannot ensure a balanced category distribution, we employ a stratified sampling algorithm, which divides the dataset into strata based on the labels. Each stratum corresponds to a distinct entity type, and we ensure an relatively equal number of samples per category by drawing from these strata, thereby maintaining balance across categories in the results. The size for each stratum is calculated with the formula:
$s_{i}=\min\left(\left\lfloor\frac{S}{m}\right\rfloor,n_{i}\right)$ (8)
where $N$ is the total number of labels in the dataset, $S$ is the total sample size, $n_{i}$ is the total number of labels with value $i$, $m$ is the number of categories, and $s_{i}$ is the number of labels from stratum $i$.
Dataset Selection
We conduct supplementary experiments on existing datasets including CoNLL2003 Tjong Kim Sang and De Meulder (2003), CrossNER Liu et al. (2021), FewNERD Ding et al. (2021), PAN-X Pan et al. (2017), and MultiCoNER Malmasi et al. (2022). Since we decide to use and share the formatted versions of these datasets in our repository to facilitate the test and use of CascadeNER, we only choose open-sourced datasets to avoid copyright issues. For evaluation metrics, we primarily use F1 score, as it is widely recognized as the most robust and effective metric for NER tasks Li and Sun (2020). We detail below the reasons for selecting these datasets and their usage.
Model CoNLL2003 AI Literature Music Politics Science FewNERD-8 FewNERD-66
XLM-RoBERTa 92.3 59.0 65.9 72.1 70.8 66.9 80.5 64.1
ACE+document-context 94.6 17.2 22.6 23.8 35.1 32.3 83.3 70.4
BERT-MRC+DSC 93.5 63.2 67.8 74.5 76.1 68.7 86.7 74.1
PromptNER 84.2 64.8 74.44 84.2 78.6 72.6 76.5 35.6
GPT-NER 73.5 58.0 61.2 60.8 62.4 55.8 70.0 58.4
CascadeNER (zero-shot) 88.2 68.9 71.7 79.3 80.5 73.6 73.4 67.0
CascadeNER (few-shot) 92.8 75.8 75.2 83.2 82.4 77.1 84.5 75.9
Table 6: F1 score of different models on CoNLL2003, CrossNER, and FewNERD.
Model PAN-X MultiCoNER
en es fr ru de zh ja ko en es ru de zh ko
XLM-RoBERTa 88.1 86.5 85.4 86.3 83.1 78.3 75.6 82.0 58.9 54.8 55.9 60.6 62.6 52.0
GEMNET 90.5 91.1 87.6 87.4 86.6 81.5 80.8 85.5 84.3 85.3 78.7 89.5 83.2 85.7
PromptNER 81.7 79.6 73.5 73.8 71.9 72.1 70.8 73.5 79.5 75.6 76.5 67.6 70.8 72.4
GPT-NER 75.2 72.8 71.6 63.5 72.0 72.4 71.5 72.1 71.7 67.9 58.2 63.1 61.2 62.5
CascadeNER (zero-shot) 87.8 85.0 83.2 80.7 77.4 78.7 74.7 72.0 71.9 71.5 71.2 63.5 70.3 69.8
CascadeNER (few-shot) 91.0 85.2 87.2 86.8 82.8 87.0 83.2 79.4 85.9 81.1 79.5 69.1 85.1 76.9
Table 7: F1 score of different models across languages on PAN-X and MultiCoNER.
CoNLL2003
CoNLL2003 is the most widely used English NER dataset, featuring four types: PER, LOC, ORG, and MISC. Supervised methods achieve excellent F1 scores of 90%-95% on this dataset. We use this dataset to compare CascadeNER and other LLM-based methods with existing supervised SOTAs in classical scenarios.
CrossNER
CrossNER is a English cross-domain dataset primarily used to evaluate a model’s cross-domain generalization and low-resource performance. It consists of five independent sub-datasets, each covering a specific domain (AI, Literature, Music, Politics, and Sciences) and containing 9-17 entity types. Since the train set for the datasets only contains 100-200 sentences, supervised methods underperform compared to LLM-based methods. We use this dataset to evaluate CascadeNER in cross-domain and low-resource scenarios.
FewNERD
FewNERD is an English dataset designed to evaluate a model’s ability to handle fine-grained entity recognition and few-shot learning, comprising 8 coarse-grained types and 66 fine-grained types. For supervised methods, FewNERD applies all 66 categories, challenging the models’ classification abilities. For few-shot methods, we use the Intra-10way setting, where the train, dev, and test sets contain non-overlapping entity types. We utilize both the 8-category and 66-category settings to evaluate CascadeNER under varying levels of classification granularity.
MultiCoNER & PAN-X
MultiCoNER and PAN-X are two widely used multilingual datasets. MultiCoNER covers 6 entity types and 11 languages, while PAN-X includes 3 entity types and 282 languages. We use 6 and 8 overlapping languages from MultiCoNER and PAN-X with DynamicNER to evaluate CascadeNER’s multilingual capabilities. It is important to note that, for the purpose of controlling variables, all methods requiring training are trained using multilingual joint training.
D.4 Experimental Results
As shown in Table 6 and 7, the results indicate that in low-resource scenarios, LLM-based methods achieve significantly better results. CascadeNER surpasses existing methods on CrossNER except Music and FewNERD, and PAN-X and MultiCoNER in some languages, achieving new SOTA performance and highlighting its exceptional generalization and capability to handle complex entity categorization. However, when handling NER tasks with ample training resources and simple classifications, LLM-based methods still lag behind existing methods, whether on the English-only CoNLL2003 or the multilingual PAN-X, indicating that supervised methods are still useful in some scenerios.
Appendix E Ablation Study
E.1 Result Fusion
In Section 4.2, we introduce our union strategy in result fusion to address the issue of extractor recall being significantly lower than precision, allowing multiple extractions for one sentence and taking the union of the results to maximize recall. For the problem of entity nesting, where different extraction rounds yield overlapping or nested entities, we adopt a length-first strategy, retaining the longer entity. Table 8 provides a example for the significantly low recall.
Dataset Precision Recall F1 Score
CoNLL2003 98.4 93.6 95.9
AI 98.7 88.0 93.1
Literature 98.3 87.8 92.7
Music 98.0 92.0 94.9
Politics 97.5 90.0 93.6
Science 98.2 85.9 91.6
Table 8: Precision, recall, and F1 Score for CoNLL2003 and CrossNER. In this experiment, both base models used in CascadeNER are Qwen2.5-7B, and the results are obtained in zero-shot scenarios.
Figure 7 presents the impact of increasing the number of extraction repetitions in zero-shot scenarios on CoNLL2003. The results show that our strategy can slightly improve recall with minimal impact on precision. Given the obvious margin effect after 3 repetitions, we ultimately select $3$ as the repetition count $k$ for other experiments. It is important to emphasize that even without repetition, CascadeNER still has a significant performance advantage.
Figure 7: The curves showing visualized precision, recall, and F1 Score as a function of the number of repetitions, demonstrating how these metrics change with increasing repetition counts $k$. Both base models used in CascadeNER are Qwen2.5-7B.
E.2 Varying the Base Models
In this section, we use four different lightweight LLMs as the base models for CascadeNER, i.e., two versions of Qwen2 and Gemma with different parameters, namely Qwen2-1.5B, Qwen2-7B, Gemma-2B, and Gemma-7B. Gemma is another prominent lightweight LLM series, proposed by Google Team et al. (2024). These models represent the current best-performing lightweight LLMs. We use the dynamic categorized version of DynamicNER to fine-tune the instruct versions of the four models. Each model is fine-tuned separately on the corresponding dataset to obtain both an extractor and a classifier. The performance comparison of these combinations on a selected part of DynamicNER is shown in Table 9. Based on these results, Qwen2.5 outperform Gemma in multilingual tasks overall. Therefore, we choose Qwen2.5 as the base models for other experiments.
English Chinese Spanish Japanese
Qwen2.5-1.5B 62.8 58.9 55.7 54.1
Qwen2.5-7B 68.2 64.5 61.5 60.8
Gemma-2B 58.9 49.5 53.1 48.4
Gemma-7B 61.7 53.9 55.3 52.1
Table 9: F1 scores for CascadeNER with different base models on a selected part of DynamicNER
E.3 Context in Classification
In the early stages of our research, the prompt used for classification contained only the entity itself without any context. Figure 8 provides an example comparing the two types of prompts. Although this method makes the prompt more concise, it lacks any contextual information. Our final in-context classification queries significantly improve classification accuracy, as shown in Table 10.
Figure 8: Example of the early context-free queries.
Dataset ACC (context-free) ACC (in-context)
CoNLL2003 90.1 94.2
AI 75.5 79.6
Literature 78.9 83.4
Music 84.6 88.3
Politics 87.4 90.8
Science 82.2 86.5
Table 10: Both base models used are Qwen2.5-7B. The results are obtained in zero-shot scenarios. The used datasets are CoNLL2003 and CrossNER. Accuracy is used in the evaluation of classifiers.
Appendix F LLM-based Methods Comparsion
In this section, we compare our prompt with two existing LLM-based baselines, GPT-NER Wang et al. (2023) and PromptNER Ashok and Lipton (2023). These methods are the currently main methods to achieve general NER with LLMs. A breif comparsion is shown in Figure 9.
Figure 9: Examples for the three parts of the prompt for each method. The red boxes contain the task description, the green boxes contain few-shot demonstrations, and the blue boxes contain the input sentence.
PromptNER utilizes detailed descriptions of each entity’s specific definition and CoT reasoning processes to fully leverage the LLM’s logical reasoning abilities. However, like traditional methods, it treats NER as a sequence labeling task, failing to effectively utilize the LLM’s global contextual understanding capabilities, making it prone to overlooking important context in complex sentences. Additionally, the task descriptions are overly complex, which not only makes it difficult for lightweight LLMs to correctly execute tasks, but also leads to a higher likelihood of hallucinations in tasks requiring fine-grained classification, such as the fine-grained settings of FewNERD and DynamicNER, as each category’s definition requires descriptions. These issues reduce PromptNER’s generalization and accuracy, limiting its application.
GPT-NER handles the NER task by determining whether a single entity belongs to a specific category, which leverages the generative capabilities of LLMs and allows for improved attention to the influence of context on entity meaning. Its drawback lies in the fact that it can only process one entity type at a time. This makes the method highly inefficient when dealing with fine-grained categorization, leading to significant resource consumption. Additionally, this method requires multiple judgments for the same entity, introducing the potential for conflicts between different rounds. Unfortunately, GPT-NER does not provide an effective solution for this issue.
CascadeNER divides the NER task into two sub-tasks: extraction and classification, while simplifying the input and output formats and reducing logical complexity. This ensures that even lightweight LLMs with limited capacity to handle complex tasks can still perform the tasks accurately and efficiently. In extraction, CascadeNER leverages the model’s generation capabilities by producing sentences with identifiers, treating all entities uniformly, which enhances the model’s generalization ability across different languages and domains. Notably, it avoids reliance on word order by consistently using "##" to mark entities, ensuring consistent annotation regardless of whether the language is right-to-left or left-to-right, improving cross-language consistency and adaptability. In classification, our method processes the entire sentence as a whole, better utilizing LLMs’ strengths in contextual understanding and semantic modeling. By leveraging the LLM’s ability to model long-range dependencies, the model’s capacity to handle complex sentence structures is enhanced, avoiding fragmentation of information and improving overall consistency and generalization. However, our method also has limitations. The use of unified identifiers prevents CascadeNER from effectively handling nested NER. We plan to address this by developing a solution that accommodates both multilingual and nested NER tasks in future.
Appendix G Computational Resource Usage Record
In Table 11, we provide the API costs incurred when testing the complete dynamic version of DynamicNER in few-shot scenarios using three LLM-based methods with GPT-4o, serving as a reference for the computational resources required by these methods. The cost calculation follows OpenAI’s official GPT-4o pricing, with input costs at 2.5 USD per 1M tokens and output costs at 10 USD per 1M tokens. The records show that CascadeNER exhibits significant advantages over existing methods in computational resource consumption.
GPT-NER PromptNER CascadeNER
Cost (USD) 513.92 128.49 45.86
Table 11: Cost comparison of three LLM-based methods. The cost is calculated according to OpenAI’s official GPT-4o pricing, not the actual cost.
Appendix H Data Contamination Statement
Given that LLMs are trained on data from diverse and complex sources, there is a possibility that portions of the evaluation sets may have been encountered during pre-training. However, as prior research Chowdhery et al. (2023) indicates, contaminated data that has been seen during training does not significantly influence performance. Thus, we consider this issue negligible.
In additional experiments on CascadeNER, we notice another critical data contamination concern: potential corpus overlap between DynamicNER and other benchmark datasets utilizing Wikipedia-derived text, which can reduce evaluation fairness. To mitigate this risk, we implement a rigorous filtering protocol during DynamicNER’s annotation phase. After completing the initial manual annotation of the base version, we employ Sentence-BERT to compute semantic cosine similarity between each candidate sentence and existing sentences in reference datasets. Sentences exhibiting similarity scores exceeding 0.8 are excluded from the corpus. New sentences from collected corpus meeting the similarity criteria are then re-annotated following the original annotation workflow. This iterative process continues until all sentences in the base version satisfy the similarity constraints. After this we utilize dynamic categorization to generates the dynamic version. This procedure ensures the reliability and fairness of our test results.
Appendix I Ethical Statement of DynamicNER
When constructing DynamicNER, we strictly adhere to existing ethical guidelines Bender and Friedman (2018); Gebru et al. (2021); Hovy and Spruit (2016), ensuring that our data sources and processing methods comply with legal and ethical standards while maintaining high-quality annotations. All the text in DynamicNER is sourced from Wikipedia, ensuring no violations of privacy or copyright, as Wikipedia is an open-source platform with user-contributed content from around the world. During data collection and annotation, we balance category distribution to minimize the risk of bias in the model. Furthermore, we maintain transparency by detailing the dataset development process and data partitioning in this paper, ensuring clarity and reproducibility for future research.
For the annotators, each language in DynamicNER is annotated by two junior or higher-level students from the corresponding language departments at our university. Due to the double-blind review process, the annotators’ identities cannot be disclosed in this version. Each annotator receives extensive training and follows DynamicNER’s multi-granularity classification system to ensure consistent and accurate entity annotations across various languages and domains. The annotation process for each language are divided into two parts equally, with each annotator independently handling one part. After the initial annotation, the annotators revise their work based on the review results. For ambiguous terms or specialized domain terms, the annotators either collaborate with each other or consult experts to ensure the accuracy and reliability of the annotations.
In our writing, we use CahtGPT-4o Achiam et al. (2023) and Claude 3.5 Anthropic (2023) for assistance.
Appendix J Detailed Categories of DynamicNER
J.1 Person
Real Person
Politician, Artist, Author, Athlete, Director, Actor, Scholar, Military, Musician, Business Executive, Other Person.
Fictional Figure
Mythological Figure, Other Figure.
J.2 Location
Geographical Entity
Water Body, Mountain, Island, Desert, Other Geographical Entity.
Geo-Political Entity
Continent, Country, State or Province, City, District, Region, Other GPE.
Address
Address, Road, Railway, Other Address.
J.3 Product
Food
Beverages, Packaged Foods, Other Food.
Weapon
Firearms, Biological, Chemical Weapon, Explosives, Cold Weapon, Nuclear, Other Weapon.
Technology
Software, Website, Electronics, AI, Other Technology.
Vehicle
Air, Car, Water, Rail, Bike, Other Vehicle.
Other Product
Clothes, Household, Personal Care, Toys, Musical Instruments, Other Product.
J.4 Facility
Public Facility
Hospital, Library, Park, Landmark, School, Museum, Sports Facility, Other Public Facility.
Commercial Facility
Hotel, Restaurant, Market/Mall, Theater/Cinema, Bank, Other Commercial Facility.
Transportation Facility
Airport, Station, Port, Other Transportation Facility.
Production Facility
Factory, Farm, Mine, Energy, Other Production Facility.
Other Facility
Residential, Government Facility, Other Facility.
J.5 Art
Visual Art
Painting, Sculpture, Visual Art Genre, Other Visual Art.
Music
Song, Album, Music Genre, Other Music.
Literature
Poem, Non-fiction, Fiction, Literature Genre, Other Literature.
Other Art
Film, Play, Broadcast Program, Game, Other Art.
J.6 Group
Social Group
Ethnic Group, Religious Group, Other Social Group.
Non-commercial Organization
Educational and Research, Political/Military, Community, Religious Organization, Other Non-commercial Organization.
Commercial Organization
Sports Team, Band, Company, Media, Other Commercial Organization.
J.7 Miscellaneous
Award
Literary Award, Sports Award, Artistic Award, Other Award.
Event
Political/Military Event, Sporting Event, Disaster, Business Event, Other Event.
Miscellaneous
Educational Degree, Tradition, God, Law, Language, Miscellaneous.
J.8 Science Entity
Biological
Protein, Species, Biological Theory, Other Biological Entity.
Chemical
Element, Compound, Reaction, Chemical Theory, Other Chemical Entity.
Physical
Physical Phenomenon, Astronomical Object, Physical Theory, Other Physical Entity.
Computer Science
ProgramLang, Algorithm, Other Computer Science Entity.
Medical
Disease, Injury, Medication, Symptom, Medical Theory, Other Medical Entity.
Other Scientific Entity
Discipline, Academic Journal, Conference, Metrics, Other Scientific Entity.
Appendix K More Categorization Quality Evaluation
In this section, we display the quantitative results of categorization metrics in Spanish, French, Russian, German, Japanese, and Korean. The results in shown in Figure 10. Experimental results demonstrate that our dynamic categorization method maintains or improves dataset quality compared to the base version in all languages.
Figure 10: Quantitative categorization metric results for 3 versions DynamicNER.